# Data Science Education in Undergraduate Physics: Lessons Learned from a Community of Practice

### v1

@@ II.1.1 Introduction to Data Processing with Histograms
This module is designed to teach students the steps involved in a data analysis pipeline. It focuses on key concepts such as processing datasets, creating histograms, curve fitting, and determining the goodness of fit for a chosen model. It covers the practical use of tools such as NumPy and SciPy for creating histograms and curve fitting.
@@
The module comprises in-class notebooks, a quiz, a homework assignment, and relevant datasets. Four notebooks guide students through the process of analyzing a dataset, including the creation of histograms and curve fitting. A short quiz is also included to assess students’ understanding. The homework assignment then requires students to undertake a practical data analysis task drawing from the material presented in the notebooks. Several datasets are provided for this purpose, along with Jupyter notebooks for instructors to re-generate the datasets if desired. It is estimated that the completion of each notebook will take approximately 60 minutes. The quiz is designed to take around 30 minutes, and the homework assignment is anticipated to require less than 90 minutes of commitment.
@@
The module explores the application of histograms as a tool for data visualization, where students are expected to create their own histogram from a raw dataset and employ built-in functions. Students are introduced to the concept of interpreting histograms as probability distributions. This concept is exemplified using a toy “Uranium-241” dataset, where students are expected to import the data, create a histogram, and subsequently normalize these histograms. The module is designed to be a part of an undergraduate laboratory course or a particle physics course.
Students are expected to achieve several physics and data science competencies after completing this module. From a physics perspective, they should demonstrate the capability to use histograms to visualize large datasets, evaluate how effectively a chosen model aligns with a dataset, and construct their own data analysis pipelines. On the data science front, the students should become adept at utilizing built-in functions from numpy and scipy to produce histograms from a dataset, using these libraries to curve fit a distribution, and even writing their own code for these tasks, if desired. A bonus advanced topic available to students is the utilization of the Kolmogorov-Smirnov goodness-of-fit test for histograms, providing an additional level of depth for those wishing to further their understanding. Pre-requisites for this module include familiarity with Python, numpy, and matplotlib, which are crucial to understanding and completing the exercises and tasks included in the module.
II.1.2 Learning the Schrödinger Equation
@@
This module aims to introduce students to deep learning as applied to the Time-Dependent Schrödinger Equation (TDSE), a fundamental equation in quantum mechanics (QM), through a combination of analytical as well as computational (ML) solutions. The module highlights Physics-Informed Neural Networks (PINNs), a class of machine learning algorithms that facilitate the solution of partial differential equations by integrating constraints from partial differential equations within deep learning models. The module is designed for an undergraduate QM course, and it guides students through solving the TDSE with a focus on the Quantum Harmonic Oscillator (QHO). The only prerequisite is an introductory understanding of Python and the basics covered in the initial weeks of a QM course.
@@
The module is divided into three submodules. The first is a gentle, interactive introduction to deep learning, where students learn to create neural network pipelines through a simple example involving RGB color mixing. The second part consists of an article that offers an overview of various ML concepts with an emphasis on physics applications. The third submodule provides an interactive notebook for solving the TDSE for a 1D QHO system, consisting of two states in superposition as they evolve in time.
@@
Supplementing the main content within the notebooks are hands-on lessons, exercises, homework and suggested projects. These materials allow students to apply what they have learned and deepen their understanding of the subject. There are code cells that students are expected to fill themselves, references for further study, take-home assignments, and suggestions for topics that build on the material and can be used as research projects. The estimated time for students to complete the module is about six hours, in addition to two hours taught by a professor in a classroom setting.
The module can be accessed via Google Colab, with GPU support for faster performance, or through a local environment using Conda. Detailed instructions are provided for both methods.
The “Learning the Schrödinger Equation” module blends traditional quantum mechanics with modern machine learning techniques. Through a structured sequence of submodules, students are exposed to the TDSE and are then led to explore it through hands-on applications of deep learning. Students are encouraged to extend this module by exploring other quantum systems, such as particle-in-a-box or double-well potential, and to apply the techniques from this module for solving partial differential equations encountered in other physical systems.
III Conclusions and Outlook
This paper highlights the imperative integration of data science into the undergraduate physics curriculum, emphasizing the growing interplay between these two fields. With a growing number of physicists employing data science tools in their research, the necessity for this incorporation becomes ever more apparent. Physics methodology is also being used to understand and extend data science tools. Physicists entering the workforce in the industry, national labs, or academia will be well served by having exposure to these methods. This mirrors the requirement for physicists to possess proficiency in mathematical methods and computer programming.
@@ III Conclusions and Outlook
We introduced DSECOP as a means to address this need. Educators interested in including data science in their courses are welcome to join the community. Educators can choose units based on their teaching subjects and their comfort with the methodology. In the future, we plan to hold workshops at national conferences where instructors can have more direct interaction with the graduate fellows who develop our modules.
@@ Acknowledgements
The DSECOP project was funded and actualized with the support of the 2021 American Physical Society Innovation Fund award. This project would not have been possible without the support of Dr. Marilena Longobardi and Dr. Wolfgang Losert. We acknowledge useful discussions with Dr. Jackson and Dr. Craig from EP3, Dr. Robert Hilborn from AAPT, and Dr. Talitha Washington from Atlanta University Center. We extend our acknowledgment to our DSECOP Fellows for their invaluable contributions, without which this project would not have achieved its success; 2022 fellows, listed in alphabetical order: Sebastian Atalla, Fatemeh Bagheri, Julie Butler, Cunwei Fan, Radha Mastandrea, and Karan Shah; and 2023 fellows, also in alphabetical order: Julie Butler, Ashley Dale, Joseph Dominicus Lap, Richard Harry, Connor Robertson, and Karan Shah. Support for WR’s participation in this education and outreach activity was provided by the CHRNS project under a partnership between the NIST and the NSF (DMR-2010792). Any mention of commercial products within this paper is for information only; it does not imply recommendation or endorsement by NIST.

### v2

@@ IV.0.1 Introduction to Data Processing with Histograms
This module, designed for an undergraduate laboratory or a particle physics course, teaches students the steps involved in a data analysis pipeline. It focuses on key concepts such as processing datasets, creating histograms, curve fitting, and determining the goodness of fit for a chosen model and covers the practical use of tools such as NumPy and SciPy for creating histograms and curve fitting.
@@
The module comprises in-class notebooks, a quiz, a homework assignment, and relevant datasets. Four notebooks guide students through the process of analyzing a dataset, including the creation of histograms and curve fitting. A short quiz assesses students’ understanding. The homework assignment then requires students to undertake a practical data analysis task drawing from the material presented in the notebooks. Several datasets are provided for this purpose, along with Jupyter notebooks for instructors to re-generate the datasets if desired. It is estimated that the completion of each notebook will take approximately 60 minutes. The quiz is designed to take 30 minutes, and the homework assignment is anticipated to require less than 90 minutes of commitment.
@@
The module explores the application of histograms as a tool for data visualization, where students are expected to create their own histogram from a raw dataset and employ built-in functions. Students are introduced to the concept of interpreting histograms as probability distributions. This concept is exemplified using a toy “Uranium-241” dataset. Students are expected to import the data, create a histogram, and normalize these histograms.
A bonus advanced topic available to students is the utilization of the Kolmogorov-Smirnov goodness-of-fit test for histograms, providing an additional level of depth for those wishing to further their understanding. Prerequisites for this module include familiarity with Python, numpy, and matplotlib.
IV.0.2 Learning the Schrödinger Equation
@@
This module aims to introduce students to deep learning as applied to the Time-Dependent Schrödinger Equation (TDSE) through a combination of analytical and machine learning methods. The module highlights physics-informed neural networks, a class of machine learning algorithms that facilitate the solution of partial differential equations by integrating constraints from partial differential equations within deep learning models. The module is designed for an undergraduate quantum mechanics course, and it guides students through solving the TDSE with a focus on the quantum harmonic oscillator. The only prerequisite is an introductory understanding of Python and the basics covered in the initial weeks of a quantum mechanics course.
@@
The module is divided into three submodules. The first is a gentle, interactive introduction to deep learning, where students learn to create neural network pipelines through a simple example involving RGB color mixing. The second part offers an overview of various ML concepts with an emphasis on physics applications. The third submodule provides an interactive notebook for solving the TDSE for a 1D quantum harmonic oscillator system for the time evolution of a superposition of two energy eigenstates.
@@
Supplementing the main content within the notebooks are hands-on lessons, exercises, homework, and suggested projects. These materials allow students to apply what they have learned and deepen their understanding of the subject. There are code cells that students are expected to fill themselves, references for further study, take-home assignments, and suggestions for extensions, such as the particle-in-a-box or the double-well potential, that build on the material and can be used as research projects. The estimated time for students to complete the module is about six hours, in addition to two hours taught by a professor in a classroom setting.
V Conclusions and Outlook
@@
In 2022, we introduced DSECOP as a means to address the need for the integration of data science into the undergraduate physics curriculum. Educators interested in including data science in their courses are welcome to join the community at https://dsecop.org. They can choose units based on their teaching subjects and their comfort with the methodology and are also welcome to engage with the community by contributing modules based on their use cases. We also hold workshops at conferences where instructors can interact more directly with the graduate fellows who develop our modules.
The DSECOP initiative represents a significant step toward modernizing physics education by embedding essential data science skills within the curriculum. This effort enhances the relevance of physics education and equips students with the tools necessary for a data-driven world. By creating a community of practice, we ensure that educators are supported and can share best practices, thereby continuously improving the integration of data science into physics education. The flexibility and accessibility of the modules developed by DSECOP are crucial for overcoming the challenges identified in our surveys, making data science education feasible even in resource-constrained environments. Through this initiative, we aim to create a robust foundation for future physicists who are well-versed in data science, thereby bridging the gap between traditional physics education and the demands of contemporary scientific research and industry.
@@ Acknowledgements
The DSECOP project was funded and actualized with the support of the 2021 American Physical Society Innovation Fund award. This project would not have been possible without the support of Dr. Marilena Longobardi and Dr. Wolfgang Losert. We acknowledge useful discussions with Dr. Jackson and Dr. Craig from EP3, Dr. Robert Hilborn from AAPT, and Dr. Talitha Washington from Atlanta University Center. We extend our acknowledgment to our DSECOP Fellows for their invaluable contributions, without which this project would not have achieved its success; 2022 fellows, listed in alphabetical order: Sebastian Atalla, Fatemeh Bagheri, Julie Butler, Cunwei Fan, Radha Mastandrea, and Karan Shah; and 2023 fellows, also in alphabetical order: Julie Butler, Ashley Dale, Joseph Dominicus Lap, Richard Harry, Connor Robertson, and Karan Shah. Support for Dr. William Ratcliff’s participation in this education and outreach activity was provided by the CHRNS project under a partnership between the NIST and the NSF (DMR-2010792). Any mention of commercial products within this paper is for information only; it does not imply recommendation or endorsement by NIST.
@@
Appendix: Survey Details
In this section, we describe additional faculty survey details. 8% of the faculty indicated they teach data science in their introductory and intermediate/advanced physics courses. Throughout this manuscript, both intermediate and advanced physics courses will be referred to as “advanced” for brevity. 2% indicated they teach data science only in introductory physics. 5% indicated that their introductory physics course that teaches data science is focused on data science (e.g., Data Science for Physicists). 19% indicated they teach data science in only intermediate or advanced physics. 12% indicated that their advanced physics course is focused on data science.
For introductory courses, data science is taught via in-class activities (N = 20), projects (N = 8), and homework exercises (N = 7). A few respondents include data science on quizzes or exams (N = 3).
The majority of respondents (N = 70) have not taught data science in their physics courses in the past 5 years. Out of the 70 respondents, 69 answered the question regarding their interest in teaching data science in their physics courses. 40.6% selected they would be interested, 40.6% were maybe, and 18.8% were not interested. Sixty-one of the respondents answered the free-response question regarding why they do not teach data science in their physics courses. These responses were thematically coded. The respondents often gave multiple reasons. The top themes that emerged are in the table 2 below.
Theme Count Example quotes
Not relevant or necessary courses taught 20 “… I have not judged it to be a priority (or even relevant) in courses I am teaching on topics such as quantum mechanics, special relativity, and ‘honors’ level mechanics.”
Challenges in changing or adding to an already full curriculum 18 “Courses are already full. Hard to decide what to drop to make room”
Faculty do not know how to use data science 13 “I don’t understand machine learning well enough to teach it. Although, it seems no one does.”
“No background in big data, ML, or AI. ”
Faculty does not know how to include data in their physics courses 13 “I don’t know how I would integrate it into courses I have taught…”
Most respondents did not see data science as relevant or necessary to the courses they teach. These respondents were not opposed to data science; some used data science and AI in their research, but they felt that teaching such skills was part of the core curriculum or could be learned elsewhere (e.g., in a computer science department). This is related to the second most identified theme, that there are challenges in changing or adding to the physics curriculum. These respondents believed that the curriculum has a lot of material to cover as is, and some were already engaging in changes. Lastly, some respondents do not have experience using data science, and others were unsure how to include it in their courses.